The information below relates to a closed economy without a government. Use the information to
answer the questions that follow:
 Autonomous consumption = R285 million;
 Autonomous investment = R300 million;
 Marginal propensity to consume = 0.4.
Q.4.1 Calculate the level of autonomous spending. (3)
Q.4.2 Calculate the size of the multiplier. (3)
Q.4.3 Calculate the equilibrium level of income
1
Expert's answer
2020-08-25T14:17:58-0400

Solution:

Q.4.1 Calculate the level of autonomous spending:

"Autonomous\; spending\;(A_{o} ) =Autonomous\; Consumption\;(C_{o}) +Autonomous\; Investment\;(I)"


"A{o} =R\;285\;Million +R\;300\;Million"


"A{o} =R\;585\;Million"


Q.4.2 Calculate the size of the multiplier:


"Multiplier=\frac{1}{(1-MPC) }"


"Multiplier=" "\frac{1}{(1-0.4) }"


"Multiplier=\frac{1}{0.6} =1.67"


Q.4.3 Calculate the equilibrium level of income:

Equilibrium:

"Y=" "\frac{1}{(1-MPC) } (C{o} +I)"


"\frac{1}{(1-0.4) } (285 +300)"


"1.67(585) =R\;976.95\;Million"

"Equilibrium \;level \;of \;income=R\;976.95\;Million"